a school administrator claims that less than 50% of the students of the school are dissatisfied by the community cafeteria service. Test this claim by using sample data obtained from a survey of 500 students of the school where 54% indicated their dissatisfaction of the community cafeteria service. use a 0.05

1
Expert's answer
2022-05-02T14:47:55-0400

The following null and alternative hypotheses for the population proportion needs to be tested:

"H_0:p\ge0.50"

"H_a:p<0.50"

This corresponds to a left-tailed test, for which a z-test for one population proportion will be used.

The z-statistic is computed as follows:


"z=\dfrac{\hat{p}-p_0}{\sqrt{\dfrac{p_0(1-p_0)}{n}}}=\dfrac{0.54-0.50}{\sqrt{\dfrac{0.50(1-0.50)}{500}}}\approx1.7889"

The p-value is "p =P(Z<1.7889)= 0.963185"

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a left-tailed test is "z_c = -1.6449."

The rejection region for this right-tailed test is "R = \{z: z<-1.6449\}."

Since it is observed that "z = 1.7889 \ge-1.6449= z_c ," it is then concluded that the null hypothesis is rejected.

Using the P-value approach: The p-value is "p =0.963185," and since "p = 0.963185 \ge 0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is not enough evidence to claim that the population proportion "p" is less than 0.50, at the "\alpha = 0.05" significance level.